Let Z = f(x,y) = 3x3 -5y2 -225x + 70y + 23. (i) Find the stationary points of z. (ii) Determine if at these points the function is at a relative maximum, relative minimum, infixion point, or saddle point. 


1
Expert's answer
2021-04-29T17:10:55-0400

Given f(x, y) =  3x3 - 5y2 - 225x + 70y + 23

(i) for stationary point: fx = 0 & fy = 0


"\implies" fx(x, y) = 9x2 - 225 = 0

9x2 = 225

x2 = "\frac{225}{9}"

x = "\sqrt{\frac{225}{9}}"

x = "\pm" "\frac{15}{3}"


"\implies" fy(x, y) = -10y + 70 = 0

-10y = -70

y = "\frac{70}{10}"

y = 7


therefore the possible stationary points are ( "\frac{15}{3}", 7) and ( -"\frac{15}{3}", 7)


(ii) now to Determine if at these points the function is at a relative maximum, relative minimum, infixion point, or saddle point. 


fxy = 0 , fxx = 18x and fyy = -10

therefore,

D = fxx"\cdot" fyy - (fxy)2 = (18x)(-10) - (0)2 = -180x


D( "\frac{15}{3}", 7) = - 180 "\cdot \frac{15}{3}" = -900 < 0

D( - "\frac{15}{3}", 7) = - 180 "\cdot \frac{-15}{3}" = 900 > 0


fxx( - "\frac{15}{3}", 7) = 18 "\cdot \frac{-15}{3}" = -90 < 0


therefore by second order extremum test,

"\implies" ( -"\frac{15}{3}", 7) is relative maximum and ( "\frac{15}{3}", 7) is saddle point